# Models of attractor dynamics in the brain

### v1

@@ Introduction
Recurrent neural networks support a class of dynamics known as attractor dynamics, where network activity evolves toward and stabilizes around specific patterns—attractors in the state space of possible neural activations (Hopfield, 1982; Amit, 1989). These attractors can manifest as point attractors (stable equilibrium states), line or ring attractors (continuous manifolds of stable states), or limit cycles (periodic trajectories), providing the substrate for persistent neural activity essential for working memory, perception, decision-making, spatial navigation, and planning. The computational richness of attractor networks stems from their ability to store multiple stable states, implement pattern completion from partial inputs, resist noise, and support information integration over time (Wang, 2001; Rolls, 2007; Tang et al., 2018)
@@ 1 Attractor dynamics in hippocampal place cells
Wills et al. (2005) investigated hippocampal attractor dynamics by examining changes in place cell firing patterns as rats explored environments of varying shapes. Rats were first familiarized over six days with two distinct enclosures, one square and one circular, during which hippocampal place cells in CA1 exhibited global remapping between the two shapes in four out of six animals (Figure 1A). On the seventh day, environments were systematically morphed from square to circle in intermediate (octagonal) shapes (square-like and circle-like morph trials were alternated), and the response of individual place cells was monitored. If the place cell firing patterns exhibit attractor dynamics, then the patterns for the square-like shapes should closely resemble the firing pattern in the square enclosure, and the circle-like shapes should closely resemble the firing pattern in the circular enclosure. Indeed, cells demonstrated an abrupt, attractor-like switch between square-like and circle-like firing patterns (Figure 1B,C), observable even within the initial seconds of each morph trial, although these transitions became increasingly pronounced with continued exposure. These findings provide strong empirical support for attractor dynamics in hippocampal spatial representations. Interestingly, contrasting results from Leutgeb et al. (2005) showed more graded transitions in similar morphing paradigms, suggesting that the hippocampus can operate in different dynamical regimes—discrete or continuous—depending on factors such as input ambiguity, learning history, and task demands (Tsodyks, 2005).
@@ 2 Attractor dynamics in inferior temporal cortex
The inferior temporal (IT) cortex is critical for visual object recognition (DiCarlo et al., 2012) and long-term visual memory storage (Sakai and Miyashita, 1991). Given its rich recurrent connectivity and involvement in memory retrieval, IT has been hypothesized to exhibit attractor dynamics similar to those observed in the hippocampus. Attractor network models have been proposed as computational mechanisms to support categorization and stabilize object representations, particularly under conditions of visual ambiguity (Miyashita et al., 1993; Rolls, 2007).
@@
To explore the underlying mechanisms, Akrami et al. (2009) implemented an autoassociative neural network model (Figure 2C) comprising two layers of neurons each: an input layer mimicking early visual representations and a recurrent output layer simulating the IT cortex. Each output neuron received sparse, randomly assigned feedforward inputs from neurons in the input layer and recurrent inputs from randomly selected neurons within the output layer.
Visual memory patterns were stored within the recurrent connections using a Hebbian covariance learning rule:
@@
Here, is the activity of neuron at time , is the summed synaptic input, is the firing threshold, is the neural gain, and parameters , , and control the strength and time-scale of adaptation.
Simulations revealed that asymmetric attractor-like convergence of neural responses emerged specifically when morph stimuli were interpolated between a stored (“effective") and an unstored (“ineffective”) pattern, and the network operated near its maximum memory storage capacity (approximately patterns in this model; Figure 2D). Under these conditions, responses to morphs resembling the stored pattern converged toward the corresponding attractor state, while responses to morphs closer to the unstored pattern remained linearly graded (i.e., dependent on morph level). In contrast, when memory load was low (e.g., patterns), convergence was overly broad with almost all morphs attracted to the stored memory, inconsistent with sharp behavioral categorization and neural responses observed empirically. Furthermore, when both morph endpoints were stored, convergence occurred symmetrically, eliminating the observed in IT responses. To explain this, the authors proposed that this discrepancy could reflect the experimental selection bias in neuron sampling: neurons were selected based on their responsiveness to one of the stimuli, likely belonging to the attractor basin of the "effective" pattern, whereas the "ineffective" pattern may have been represented elsewhere in IT (Haxby et al., 2001; Kiani et al., 2007). Thus, the observed asymmetry likely reflects this experimental bias in selective sampling of neuronal populations belonging predominantly to one endpoint’s memory representation rather than a fundamental computational difference between stored and unstored patterns. Finally, the inclusion of spike-frequency adaptation improved the model’s match to experimental data by allowing network activity to decay over time after stimulus offset and effectively replicating the experimentally observed temporal response dynamics.
@@ 3 Prior experience and perceptual biases
Perception is not a passive reflection of sensory input, but a constructive process profoundly shaped by prior experience that gives rise to systematic perceptual biases (Helmholtz, 1924; Beck, 1967). When faced with ambiguous stimuli, the brain actively resolves uncertainty by integrating contextual information and drawing on memory-based predictions (Friston, 2005; Bar, 2007). Prior exposure to prototypical stimuli can bias perception in two seemingly opposing directions: adaptation aftereffects, where perception is repelled away from the recently experienced adapter stimulus, and priming effects, where perception is attracted toward it (Logothetis and Pauls, 1995; Tulving and Schacter, 1990). A compelling example occurs in facial emotion perception, where prolonged exposure to a happy face causes a subsequent neutral face to appear slightly angry—a classic repulsive aftereffect (Webster et al., 2004; Aguado et al., 2007), as shown in Figure 3. However, a brief exposure to a happy face can cause subsequent faces to be perceived happier-a classic priming effect (Murphy and Zajonc, 1993).
@@
In a different take, Akrami et al. (2010) offered a unified mechanistic explanation using attractor neural network models incorporating firing rate adaptation. In this framework, the same visual input can produce either adaptation or priming depending on the temporal dynamics of network activity and the stability of stored memory representations. Specifically, short-lived inputs may nudge activity toward a familiar attractor, yielding priming, while sustained stimulation can destabilize that attractor via adaptation, producing repulsion. This model provides an account of how opposing perceptual biases emerge from a single underlying circuit architecture modulated by memory and experience (see below for details).
@@
The authors conducted three behavioral experiments using morphed images of animals, plants, and objects under varying temporal conditions (Figure 4A). In the first experiment, participants were presented with a clear prototype image (adapter) followed shortly (ms delay) by a morph between that prototype and another object. Participants consistently exhibited a repulsive bias, perceiving the morph more dissimilar to the adapter than it actually was (Figure 5-top), demonstrating adaptation aftereffects in the perception of complex, non-face objects. To establish the temporal stability of these adaptation aftereffects, the second experiment introduced a longer delay between adapter and target (ms). Surprisingly, the repulsive effect disappeared and was replaced by an attractive priming effect—participants were now more likely to judge the ambiguous morph as resembling the adapter (Figure 5-bottom). This temporal reversal suggests that adaptation effects weaken over time, revealing a slower-developing priming mechanism that biases perception toward previously seen stimuli. In the final experiment, the researchers tested whether adaptation could still occur when the adapter itself was ambiguous. Unlike in previous experiments, adaptation to an ambiguous stimulus consistently led to priming, regardless of the duration of the delay. This finding suggests that adaptation requires a well-defined, strongly encoded prototype to exert a repulsive influence on perception, while ambiguous stimuli are more likely to produce attractive biases.
Collectively, these findings demonstrate that the same stimuli and task can elicit either adaptation or priming effects, depending on temporal task parameters, highlighting the dynamic nature of perceptual biases. The results point to two opposing yet interrelated neural mechanisms: adaptation, which induces repulsive aftereffects by attenuating responses to recently encountered stimuli, and memory-based priming, which attracts perception toward prior stimuli via the activation of stable memory representations. When a stimulus is well-defined and recently presented, adaptation dominates, pushing perception away. As time passes, or when adaptation is weak (as with ambiguous stimuli), priming emerges, pulling perception toward familiar experiences. This work offers a compelling demonstration of how perceptual history modulates ongoing experience and provides key insights into the temporal dynamics that govern the balance between adaptation and memory-based priming.
@@
The studies reviewed above showed that perception is inherently contextual, relying on the integration of prior experience with incoming sensory input to dynamically resolve categorical decisions. This integration is fundamentally rooted in the temporal dynamics of neurons involved in perceptual computations. To address this, Akrami et al. (2010) developed a neural network model that unifies historically separate perceptual phenomena-adaptation aftereffects and priming-within a single theoretical framework grounded in attractor dynamics. Their model reproduces key electrophysiological and behavioral findings from Daelli et al. (2010) and accounts for a range of priming effects observed in humans.
@@
The network architecture is an auto-associative memory model comprising recurrently connected neurons with threshold-linear activation functions. Crucially, it incorporates firing rate adaptation, a biophysical property of pyramidal neurons, which allows the network to transition between transient and stable states. The investigation centers on the interplay between recent input-driven activity and stable attractor dynamics to show how this interplay gives rise to systematic perceptual biases.
@@
To quantify these effects, the model defines an index measuring the distance between two network states at any time:
@@
where is the overlap with pattern A when A is the adapter, and is the overlap with A when B is the adapter. This metric allowed the authors to systematically map the influence of adaptation strength, delay, and target duration on perceptual outcomes across a broad range of parameter space.
The simulations revealed that firing rate adaptation is the essential component for producing adaptation aftereffects (Figure 6B), while two distinct mechanisms underlying priming effects were identified:
Perceptual Bias (Type A Priming): When the adapter induces an attractor state and the target provides a weak or ambiguous input, the network remains in the prior state, biasing perception toward the adapter. This occurs primarily when adaptation is minimal (Figure 6A).
@@
Double Adaptation (Type B Priming): Under stronger firing rate adaptation, the network can undergo multiple transitions between attractor states. Under specific temporal conditions, this results in the network settling back into the attractor associated with the adapter, paradoxically producing a priming effect even in the presence of adaptation. (Figure 6C).
@@
The research also examined backward masking effects, where a secondary stimulus disrupts the processing of the initial adapter. Simulations revealed that if a mask is presented before the network fully settles into an attractor state, adaptation aftereffects are substantially weakened, leaving predominantly priming effects. This provides a mechanistic explanation for previously observed interactions between masking and adaptation.
By systematically varying parameters—including memory load (i.e., number of stored patterns), firing rate adaptation strength, and the nature of adapter stimulus (ranging from well-defined prototypes to ambiguous morphs)—the model maps how network dynamics shift between adaptation aftereffects and priming regimes. Robust adaptation effects were shown to require strong firing rate adaptation, while priming emerged in networks with stable attractor representations, contingent on the timing of the adapter and target presentation.
The temporal relationship between stimuli proved critical. Brief adapter-target intervals favored adaptation aftereffects, while extended delays promoted priming. Moreover, target duration emerged as an additional factor modulating outcomes: longer target exposures could reverse expected effects by allowing the network to overcome adaptation and converge on a stored attractor. The model’s predictions align closely with behavioral findings from face and motion adaptation studies, reinforcing the model’s relevance to real-world perceptual dynamics.
These findings underscore the importance of firing rate adaptation as a key modulator of perceptual dynamics and highlight how attractor networks provide a mechanistic basis for the flexible, experience-dependent biases observed in human perception.
@@ 4 Coupled PPC and PFC Dynamics Underlying Biases in Working Memory
A key aspect of the PWM task is the manipulation of stimulus pairs (presented at each trial) based on their proximity to the identity line = (Figure Figure 7B), which changes the comparison difficulty. Pairs closer to the diagonal line are harder to discriminate due to smaller intensity differences. Rats were presented with stimulus pairs that were all equally distant from the diagonal (see Figure 7B). Despite the equal objective difficulty, performance was systematically modulated by contraction bias: the remembered value of appeared to be shifted toward the average of previously encountered stimuli (see Figure 7A). When this shift increased the perceived distance from (bias+), performance improved; when it reduced the difference (bias-), performance declined.
To quantify these history affects, Akrami et al. (2018) used a logistic regression to model behavioral choices as a function of current stimulus values , the average of past stimuli (capturing contraction bias), the stimulus pair from the immediately preceding trial (capturing recency), and reward and choice history. This analysis revealed clear evidence for both contraction and recency biases. Importantly, similar biases were observed in human participants, suggesting conserved underlying mechanisms.
@@
To complement the mechanistic account, Boboeva et al. (2024) also provided a probabilistic model, summarizing the network dynamics. In this simpler framework, the memory of either remains veridical (unchanged with high probability) or is resampled from a distribution reflecting past stimuli. Despite its simplicity, this model captured key behavioral patterns and mirrored predictions of the neural network model, providing convergent evidence that history-dependent biases in working memory can emerge from probabilistic inference shaped by PPC-WM neural dynamics.

### v2

@@ Introduction
Recurrent neural networks support a class of dynamics known as attractor dynamics, where network activity evolves toward and stabilizes around specific patterns—attractors in the state space of possible neural activations (Hopfield, 1982; Amit, 1989). These attractors can manifest as point attractors (stable equilibrium states), line or ring attractors (continuous manifolds of stable states), or limit cycles (periodic trajectories), providing the substrate for persistent neural activity essential for working memory, perception, decision-making, spatial navigation, and planning. The computational richness of attractor networks stems from their ability to store multiple stable states, implement pattern completion from partial inputs, resist noise, and support information integration over time (Wang, 2001; Rolls, 2007; Tang et al., 2018).
@@
Broader landscape. Here we focus on discrete (point) attractors and autoassociative networks because they map cleanly onto the examples covered in the lectures. Attractor theory spans a wider set of influential models. For example, continuous (“bump") attractors have been central for parametric working memory and spatial representations (Compte et al., 2000; Wimmer et al., 2014). Line-attractor constructions provide a limiting case for stable integration and graded memory (Seung, 1996). In decision-making, recurrent attractor networks offer a mechanistic account of categorical choices and speed–accuracy trade-offs (Wang, 2002; Wong and Wang, 2006). These related formulations share the core idea of low-dimensional stable manifolds in population activity, but differ in what is stored (discrete states vs. continuous variables) and which behavioral signatures they are designed to explain.
@@ 1 Attractor dynamics in hippocampal place cells
Wills et al. (2005) investigated hippocampal attractor dynamics by examining changes in place cell firing patterns as rats explored environments of varying shapes. Rats were first familiarized over six days with two distinct enclosures, one square and one circular, during which hippocampal place cells in CA1 exhibited global remapping between the two shapes in four out of six animals (Figure 1A; the remaining two animals either did not show remapping between square and circle enclosures or their remapping did not generalize across enclosures material despite having the same shape). On the seventh day, environments were systematically morphed from square to circle in intermediate (octagonal) shapes (square-like and circle-like morph trials were alternated), and the response of individual place cells was monitored. If the place cell firing patterns are driven by a latent attractor network, then the patterns for the square-like shapes should closely resemble the firing pattern in the square enclosure, and the circle-like shapes should closely resemble the firing pattern in the circular enclosure. Indeed, cells demonstrated an abrupt, attractor-like switch between square-like and circle-like firing patterns (Figure 1B,C), observable even within the initial seconds of each morph trial, although these transitions became increasingly pronounced with continued exposure. These findings provide strong empirical support for attractor dynamics in hippocampal spatial representations. Interestingly, contrasting results from Leutgeb et al. (2005) showed more graded transitions in similar morphing paradigms, suggesting that the hippocampus can operate in different dynamical regimes—discrete or continuous—depending on factors such as input ambiguity, learning history, and task demands (Tsodyks, 2005). Consistent with these observations, computational implementations of hippocampal attractor dynamics in morphing-environment paradigms show that recurrent network dynamics can settle into one of the previously learned spatial maps (often with hysteresis) despite ambiguous intermediate inputs, producing attractor-like transitions between map states (Rennó-Costa et al., 2014; Romani and Tsodyks, 2010).
@@ 2 Attractor dynamics in inferior temporal cortex
The inferior temporal (IT) cortex is critical for visual object recognition (DiCarlo et al., 2012) and long-term visual memory storage (Sakai and Miyashita, 1991). Given its rich recurrent connectivity and involvement in memory retrieval, IT has been hypothesized to exhibit attractor dynamics similar to those observed in the hippocampus. Attractor network models have been proposed as computational mechanisms to support categorization and stabilize object representations, particularly under conditions of visual ambiguity (Miyashita and others, 1993; Rolls, 2007).
@@
To explore the underlying mechanisms, Akrami et al. (2009) implemented an autoassociative neural network model (Figure 2C) comprising two layers of neurons each: an input layer mimicking early visual representations and a recurrent output layer simulating the IT cortex. Each output neuron received sparse, randomly assigned feedforward inputs from neurons in the input layer (modeling neurons in early/mid visual cortex that exhibit sparse activations, Willmore et al. (2011)) and recurrent inputs from randomly selected neurons within the output layer.
Visual memory patterns were stored within the recurrent connections using a standard Hebbian covariance learning rule (Rolls and Treves, 1997):
@@
Here, is the activity of neuron at time , is the summed synaptic input, is the firing threshold, is the neural gain (characterizes the response proportion of a neuron given its input drive), corresponds to rectified linear activation (ReLU), and parameters , , and control the strength and time-scale of adaptation.
Simulations revealed that asymmetric attractor-like convergence of neural responses emerged specifically when morph stimuli were interpolated between a stored (“effective") and an unstored (“ineffective”) pattern, and the network operated near its maximum memory storage capacity (approximately patterns in this model; Figure 2D). Under these conditions, responses to morphs resembling the stored pattern converged toward the corresponding attractor state, while responses to morphs closer to the unstored pattern remained linearly graded (i.e., dependent on morph level). In contrast, when memory load was low (e.g., patterns), convergence was overly broad with almost all morphs attracted to the stored memory, inconsistent with sharp behavioral categorization and neural responses observed empirically. Furthermore, when both morph endpoints were stored, convergence occurred symmetrically, eliminating the observed asymmetry in IT responses. To explain this, the authors proposed that this discrepancy could reflect the experimental selection bias in neuron sampling: neurons were selected based on their responsiveness to one of the stimuli, likely belonging to the attractor basin of the "effective" pattern, whereas the "ineffective" pattern may have been represented elsewhere in IT (Haxby et al., 2001; Kiani et al., 2007). Thus, the observed asymmetry likely reflects this experimental bias in selective sampling of neuronal populations belonging predominantly to one endpoint’s memory representation rather than a fundamental computational difference between stored and unstored patterns. Finally, the inclusion of spike-frequency adaptation improved the model’s match to experimental data by allowing network activity to decay over time after stimulus offset and effectively replicating the experimentally observed temporal response dynamics.
@@ 3 Prior experience and perceptual biases
Perception is not a passive reflection of sensory input, but a constructive process profoundly shaped by prior experience that gives rise to systematic perceptual biases (Helmholtz, 1924; Beck, 1967). When faced with ambiguous stimuli, the brain actively resolves uncertainty by integrating contextual information and drawing on memory-based predictions (Friston, 2005; Bar, 2007). Prior exposure to prototypical stimuli can bias perception in two seemingly opposing directions: adaptation aftereffects, where perception is repelled away from the recently experienced adapter stimulus, and priming effects, where perception is attracted toward it (Logothetis and Pauls, 1995; Tulving and Schacter, 1990). A compelling example occurs in facial emotion perception, where prolonged exposure to a happy face causes a subsequent neutral face to appear slightly angry—a classic repulsive aftereffect (Webster et al., 2004; Aguado et al., 2007), as shown in Figure 3. However, a brief exposure to a happy face can cause subsequent faces to be perceived happier—a classic priming effect (Murphy and Zajonc, 1993).
@@
In a different take, Akrami and others (2010) offered a unified mechanistic explanation using attractor neural network models incorporating firing rate adaptation. In this framework, the same visual input can produce either adaptation or priming depending on the temporal dynamics of network activity and the stability of stored memory representations. Specifically, short-lived inputs may nudge activity toward a familiar attractor, yielding priming, while sustained stimulation can destabilize that attractor via adaptation, producing repulsion. This model provides an account of how opposing perceptual biases emerge from a single underlying circuit architecture modulated by memory and experience (see below for details).
@@
The authors conducted three behavioral experiments using morphed images of animals, plants, and objects under varying temporal conditions (Figure 4A). In the first experiment, participants were presented with a clear prototype image (adapter) followed shortly (ms delay) by a morph between that prototype and another object. Participants consistently exhibited a repulsive bias, perceiving the morph more dissimilar to the adapter than it actually was when prompted to choose which of the two original images the morph stimulus was closest to (Figure 5-top), demonstrating adaptation aftereffects in the perception of complex, non-face objects. To establish the temporal stability of these adaptation aftereffects, the second experiment introduced a longer delay between adapter and target (ms). Surprisingly, the repulsive effect disappeared and was replaced by an attractive priming effect—participants were now more likely to judge the ambiguous morph as resembling the adapter (Figure 5-bottom). This temporal reversal suggests that adaptation effects weaken over time, revealing a slower-developing priming mechanism that biases perception toward previously seen stimuli. In the final experiment, the researchers tested whether adaptation could still occur when the adapter itself was ambiguous. Unlike in previous experiments, adaptation to an ambiguous stimulus consistently led to priming, regardless of the duration of the delay. This finding suggests that adaptation requires a well-defined, strongly encoded prototype to exert a repulsive influence on perception, while ambiguous stimuli are more likely to produce attractive biases.
Collectively, these findings demonstrate that the same stimuli and task can elicit either adaptation or priming effects, depending on temporal task parameters, highlighting the dynamic nature of perceptual biases. The experiments reveal two opposing yet interrelated effects: adaptation, which can induce repulsive aftereffects by attenuating responses to recently encountered stimuli, and memory-based priming, which attracts perception toward prior stimuli via the activation of stable memory representations. When a stimulus is well-defined and recently presented, adaptation dominates, pushing perception away. As time passes, or when adaptation is weak (as with ambiguous stimuli), priming emerges, pulling perception toward familiar experiences. Together, these results provide a clear empirical characterization of how perceptual history modulates ongoing experience and set the stage for the following section, where we show these effects can be captured mechanistically within a unified attractor network framework, governed by the balance between adaptation and memory-based priming.
@@
The studies reviewed above demonstrate that perception is inherently contextual, requiring the integration of prior experience with incoming sensory input to resolve categorical decisions. This process is fundamentally rooted in the temporal dynamics of the underlying neural circuitry. To investigate this, Akrami and others (2010) sought to bridge the gap between historically disparate perceptual phenomena—adaptation aftereffects and priming—by developing a unified neural network model grounded in attractor dynamics. The authors’ primary intention was to demonstrate that these opposing perceptual biases do not require separate mechanisms but instead emerge from the intrinsic interaction between recurrent excitation and firing rate adaptation. By situating these effects within a single biophysical framework, their model successfully reproduces key electrophysiological and behavioral findings from Daelli et al. (2010) and provides a mechanistic account for the diverse range of priming effects observed in humans.
@@
The network architecture is an auto-associative memory model comprising recurrently connected neurons with threshold-linear activation functions. Crucially, it incorporates firing rate adaptation, a biophysical property of pyramidal neurons, which allows the network to transition between transient and stable states. The investigation centers on how the interplay between recent input-driven activity and stable attractor dynamics gives rise to systematic perceptual biases.
@@
To quantify these effects in the model given different stimuli and under different experimental settings, the authors define a similarity index measuring the distance between two network states at any time:
@@
where is the overlap with pattern A when A is the adapter, and is the overlap with A when B is the adapter. This metric allowed the authors to systematically map the influence of adaptation strength, delay, and target duration on perceptual outcomes across a broad range of parameter space and as a function of time in a given trial.
The simulations reveal that firing rate adaptation (FRA) is the essential component for producing adaptation aftereffects, acting as a time-dependent negative feedback mechanism that modulates the stability of attractor states. By systematically varying parameters—including memory load, FRA strength, and stimulus ambiguity—the model maps how network dynamics shift between distinct perceptual regimes (Figure 6).
Perceptual Bias (Type A Priming): In regimes where adaptation is minimal or absent, dynamics are dominated by recurrent collateral excitation. Synaptic reverberation allows the network to remain in the induced attractor state even after the adapter is removed. Consequently, the network remains biased toward the original attractor upon target presentation (Figure 6a).
Adaptation Aftereffects (Negative Bias): Under strong firing rate adaptation, the sustained firing of neurons representing the adapter pattern builds up an adaptation current (). This increases the firing threshold for those specific neurons, effectively destabilizing the active attractor. When an ambiguous target is presented, the network is “repelled” from the fatigued state and settles into a competing attractor (Figure 6b).
@@
Double Adaptation (Type B Priming): Under stronger firing rate adaptation, the network may undergo multiple transitions between attractor states. If the initial adaptation is strong enough to push the network out of the first attractor during a delay period, specific temporal conditions allow the network to settle back into the original attractor upon target presentation, paradoxically producing priming (Figure 6c).
@@
The research also examined backward masking effects, where a secondary stimulus disrupts initial processing. If a mask is presented before the network settles into an attractor, the build-up of the adaptation current is insufficient to destabilize the state, substantially weakening aftereffects and leaving predominantly priming effects. Furthermore, the temporal relationship between stimuli proved critical. Brief adapter-target intervals favored adaptation aftereffects due to peak levels, while extended delays or longer target exposures promoted priming as the adaptation current decayed, allowing stable attractor representations to regain dominance.
By integrating these diverse factors into a unified framework, the model demonstrates that the complex interplay between recurrent excitation and firing rate adaptation provides a robust neural mechanism capable of reproducing the full spectrum of observed behaviors. It reveals that shifting experimental conditions—such as stimulus timing, masking, or ambiguity—simply move the system across different operational regimes of the same underlying attractor landscape. Ultimately, these findings underscore the importance of firing rate adaptation as a key modulator of perceptual dynamics and highlight how attractor networks provide a mechanistic basis for the flexible, experience-dependent biases observed in human perception.
@@ 4 Coupled PPC and PFC Dynamics Underlying Biases in Working Memory
A key aspect of the PWM task is the systematic manipulation of stimulus pairs (presented at each trial) based on their sensory values and proximity to the identity line (Figure 7B). Pairs closer to the diagonal line differ less in intensity. Rats were presented with stimulus pairs that were all equally distant from the diagonal (see Figure 7B). Despite this uniform structure, performance was systematically modulated by contraction bias: the remembered value of appeared to be shifted toward the average of previously encountered stimuli (see Figure 7A). When this shift increased the perceived distance from (bias+), performance improved; when it reduced the difference (bias-), performance declined.
To quantify these history effects, Akrami et al. (2018) used a logistic regression to model behavioral choices as a function of current stimulus values , the average of past stimuli (capturing contraction bias), the stimulus pair from the immediately preceding trial (capturing recency), and reward and choice history. This analysis revealed clear evidence for both contraction and recency biases. Importantly, similar biases were observed in human participants, suggesting conserved underlying mechanisms.
@@
To reveal the essential statistical mechanism underlying contraction bias in the attractor network, Boboeva et al. (2024) also formulated a simplified probabilistic model that abstracts the network dynamics. In this framework, the comparison between the remembered first stimulus and the second stimulus is treated probabilistically: on a fraction of trials, the memory of remains veridical, while on others it is replaced by a sample drawn from the distribution of PPC bump locations, reflecting accumulated sensory history. Behavioral errors arise when this sampled value leads to an incorrect comparison with , and the probability of such errors can be computed analytically as integrals over the sampling distribution. Fitting this model to both the network output and empirical behavior of rats and humans shows that it captures key signatures of contraction bias with only a single free parameter, demonstrating that history-dependent biases in the comparison can emerge as a statistical consequence of PPC-driven sampling rather than gradual drift of the memory trace. This probabilistic abstraction thus provides a compact, predictive link between mechanistic network dynamics and observed behavior, and can be related to normative Bayesian descriptions under specific assumptions.